OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−4, −1) and (4, 5)? Numerical Answers Expected!

OpenStudy (anonymous):

if you use \[\frac{ y_{1}-y_{2} }{ x_{1}-x_{2} }\] to find the slope of the line that passes by then you could plug that into \[y - y_{1}=m(x-x_{1})\] and solve for y then you will have your \[y = mx + b\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem :)